# Riemann-Silberstein geometric phase for high-dimensional light manipulation

### v1

@@ Results
The RS phase can emerge in light transmission and reflection at an interface due to the evolution of 4D polarization. Such polarization evolution is determined by the eigen properties of the media separated by the interface. As shown in Fig. 2a, we consider that a plane wave propagates along +z+z direction and normally impinges on the surface of an isotropic medium with permittivity $\varepsilon$ and permeability $\mu$. The incident electric (magnetic) field is linearly polarized in $x$ ($y$) direction. We denote the RS polarization states of the incident, reflected, and transmitted waves as $\left|F_{1}^{\text{in}}\right\rangle,\left|F_{1}^{\text{ref}}\right\rangle$, and $\left|F_{1}^{\text{tra}}\right\rangle$, respectively. The transmission involves the polarization evolution $\left|F_{1}^{\text{in}}\right\rangle\rightarrow\left|F_{1}^{\text{tra}}\right\rangle$, while the reflection involves the polarization evolution $\left|F_{1}^{\text{in}}\right\rangle\rightarrow\left|F_{1}^{\text{tra}}\right\rangle\rightarrow\left|F_{1}^{\text{ref}}\right\rangle$. These RS polarization evolutions induce the RS phases $\Phi_{\mathrm{rs}}^{\text{tra}}$ and $\Phi_{\mathrm{rs}}^{\mathrm{ref}}$, which manifest in the transmitted and reflected waves, respectively. As an example, we assume the medium is a lossless metal. In this case, $\left|F_{1}^{\mathrm{in}}\right\rangle=\frac{1}{\sqrt{2}}(\hat{\mathbf{e}}_{x}+\mathrm{i}\hat{\mathbf{h}}_{y}),\left|F_{1}^{\mathrm{ref}}\right\rangle=\frac{1}{\sqrt{2}}(\hat{\mathbf{e}}_{x}-\mathrm{i}\hat{\mathbf{h}}_{y})$, and $\left|F_{1}^{\text{tra}}\right\rangle=\frac{\sqrt{\left|{\mu}\right|}}{\sqrt{\left|{\mu}\right|+\left|{\varepsilon}\right|}}(\hat{\mathbf{e}}_{x}+\frac{\mathrm{i\sqrt{\varepsilon}}}{\sqrt{\mu}}\hat{\mathbf{h}}_{y})$. Notably, $\left|F_{1}^{\text{tra}}\right\rangle$ is linearly polarized in the constitutive frame with orientation angle $\alpha_{\mathrm{rs}}=\text{arctan}(\frac{\mathrm{i\sqrt{\varepsilon}}}{\sqrt{\mu}})$, which locates on the equator of the RS-sphere with azimuthal angle $\phi_{\mathrm{rs}}=2\alpha_{\mathrm{rs}}$. Thus, changing the material properties $(\varepsilon,\mu)$ leads to rotation of local constitutive frame and thus the variation of $\phi_{\mathrm{rs}}$. Figure 2b shows the polarization evolution paths on the RS-sphere for the transmission and reflection. We calculate the RS phases $\Phi_{\mathrm{rs}}^{\text{tra}}$ and $\Phi_{\mathrm{rs}}^{\text{ref}}$ by evaluating the solid angles $\Omega$ enclosed by the paths. The results are shown as the circles in Fig. 2c for different orientation angle $\alpha_{\mathrm{rs}}$, which agree with the transmission and reflection phases (solid lines) predicted by the Fresnel’s equation [56]. This relationship holds for normal incidence at general interfaces between isotropic media (Supplementary Note III).
Both the PB and RS phases can emerge at an interface involving anisotropic media. We consider the interface between air and a non-magnetic medium that has anisotropic in-plane permittivity $\overleftrightarrow{\varepsilon}$ with elements $\varepsilon_{xx}=\varepsilon+\delta\cos\phi,\varepsilon_{yy}=\varepsilon-\delta\cos\phi,\varepsilon_{xy}=\delta\sin\phi e^{-\mathrm{i}\theta}$, and $\varepsilon_{yx}=\delta\sin\phi e^{\mathrm{i}\theta}$. The medium supports two orthogonal 4D eigen polarization states. As shown in Fig. 2d, under the normal incidence of a plane wave with polarization $\left|\Psi_{\text{in}}\right\rangle=\frac{1}{2}[(\hat{\mathbf{e}}_{x}+\mathrm{i}\hat{\mathbf{e}}_{y})+\mathrm{i}(\hat{\mathbf{h}}_{y}-\mathrm{i}\hat{\mathbf{h}}_{x})]$, two transmitted waves $\left|\Psi_{\text{tra}}^{\pm}\right\rangle=\frac{1}{\sqrt{1+|\varepsilon\pm\delta|}}(\mathbf{E}_{\pm}+\mathrm{i}\sqrt{\varepsilon\pm\delta}\mathbf{H}_{\pm})$ and two reflected waves $\left|\Psi_{\text{ref}}^{\pm}\right\rangle=\frac{1}{\sqrt{2}}\left(\mathbf{E}_{\pm}-\mathrm{i}\mathbf{H}_{\pm}\right)$ emerge, where $\mathbf{E}_{+}=\cos\frac{\phi}{2}\hat{\mathbf{e}}_{x}+\sin\frac{\phi}{2}e^{\mathrm{i}\theta}\hat{\mathbf{e}}_{y},\mathbf{E}_{-}=\sin\frac{\phi}{2}e^{-\mathrm{i}\theta}\hat{\mathbf{e}}_{x}-\cos\frac{\phi}{2}\hat{\mathbf{e}}_{y},\mathbf{H}_{+}=\cos\frac{\phi}{2}\hat{\mathbf{h}}_{y}-\sin\frac{\phi}{2}e^{\mathrm{i}\theta}\hat{\mathbf{h}}_{x}$, and $\mathbf{H}_{-}=\sin\frac{\phi}{2}e^{-\mathrm{i}\theta}\hat{\mathbf{h}}_{y}+\cos\frac{\phi}{2}\hat{\mathbf{h}}_{x}$. Correspondingly, four polarization evolutions occur at the interface: $\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{+}\right\rangle,\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{-}\right\rangle,\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{+}\right\rangle\rightarrow\left|\Psi_{\text{ref}}^{+}\right\rangle$, and $\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{-}\right\rangle\rightarrow\left|\Psi_{\text{ref}}^{-}\right\rangle$. As an example, Fig. 2e depicts the polarization evolutions associated with $\left|\Psi_{\text{tra}}^{+}\right\rangle$ and $\left|\Psi_{\text{ref}}^{+}\right\rangle$ for $\varepsilon=1+\mathrm{i},\delta=2,\phi=30^{\circ}$, and $\theta=0^{\circ}$. Notably, the electric and magnetic polarizations evolve along the same pathway on the E-sphere and H-sphere, which are independent of the RS polarization evolution on the RS-sphere. Figure 2f shows the total geometric phases (i.e., sum of PB and RS phases) induced by the four polarization evolutions for different $\alpha_{\mathrm{e}}=\frac{\phi}{2}$. The geometric phases (circles) agree with the transmission and reflection phases (solid lines) predicted by the Fresnel’s equations. Thus, the well-known reflection and transmission phases at interfaces can be interpretated as the geometric phases induced by 4D polarization evolution.
@@
Figure 3: Wave deflection by the RS metasurface. (a) Schematic of the RS metasurface under the normally incident plane wave with linearly polarized electric field. The inset shows the structure of the meta-atom. (b) RS polarization evolution on the RS-sphere induced by the metasurface. (c) Theoretical RS geometric phase (symbols), simulated phase (solid lines), and simulated amplitude for the forward and backward scattering fields as a function of the orientation angle $\alpha_{\mathrm{rs}}$ of the RS dipole. (d) Simulated electric field scattered by the RS metasurface. (e) Experiment setup and the fabricated metasurface prototype. (f) Simulated and experimentally measured far-field intensity pattern. The shaded area marks the measurement blind zone [-15 deg, 15 deg] due to the source antenna obstruction.
RS metasurfaces and experiments
@@
Figure 3b shows the RS polarization evolution in the channel $|-1\rangle\rightarrow|+1\rangle$. The incident state $\left|F_{1}^{\text{in}}\right\rangle$ and backward scattering state $\left|F_{1}^{\text{out}}\right\rangle$ are represented by the south and north poles, respectively. The polarization state of the RS dipole $\mathbf{D}_{\mathrm{rs}}$, denoted as $\left|D_{\mathrm{rs}}\right\rangle$, is located on the equator. The orientation angle of $\mathbf{D}_{\mathrm{rs}}$ in the constitutive frame is $\alpha_{\mathrm{rs}}=\text{arctan}(\frac{\mathrm{i}m_{y}}{p_{x}})$. Rotating the meta-atoms around $x$ or $y$ direction will alter $\alpha_{\mathrm{rs}}$ (Supplementary Note IV) and change the azimuthal angle on the RS-sphere. The resulting RS phase is $\Phi_{\mathrm{rs}}=-2\alpha_{\mathrm{rs}}$. Notably, the PB phase vanishes in this case because the meta-atom rotation around $x$ or $y$ direction does not change the electric or magnetic polarization. Figure 3c shows the phases and amplitudes of the scattering fields as a function of the orientation angle $\alpha_{\mathrm{rs}}$ of the RS dipole $\mathbf{D}_{\mathrm{rs}}$. The simulated phases (solid lines) agree with the RS phase (circles) obtained by evaluating the solid angle in the RS-sphere. Notably, the phase of the backward scattering field $|+1\rangle$ exhibits a linear relationship with $\alpha_{\mathrm{rs}}$, while the phase of the forward scattering field $|-1\rangle$ is independent of $\alpha_{\mathrm{rs}}$. In addition, the scattering amplitudes remain approximately constant for different $\alpha_{\mathrm{rs}}$. These results indicate that the RS metasurface can enable Janus-type wavefront shaping and functionality.
@@
Figure 4: Wave deflection by the higher-order RS metasurface. (a) Schematic of the quadruplex RS meta-deflector under the normal incidence of a plane wave with LCP electric field. (b) 4D polarization evolution on the Poincaré hypersphere induced by the higher-order RS metasurface. (c) Theoretical geometric phase (symbols), simulated phase (solid lines), and simulated amplitude in four polarization evolution channels as a function of the orientation angle of RS dipole in the meta-atom. (d) Simulated electric field distribution of four output waves. (e) Fabricated metasurface prototype. (f) Simulated and experimentally measured far-field intensity pattern under orthogonal incidence. The shaded area marks the measurement blind zone [-15 deg, 15 deg] due to the source antenna obstruction.
@@
The RS metasurface in Fig. 3a can generate RS phase gradient along $y$ direction, which can deflect the incident wave. Figure 3d shows the simulated electric field $E_{x}$ scattered by the RS metasurface. We note that only the backward scattering field is deflected into oblique direction, and the wavefront is consistent with the deflection angle $\beta=\arcsin(\frac{1}{k_{0}}\frac{\Delta\Phi_{\mathrm{rs}}}{\Delta y})$ predicted by the generalized Snell’s law [57]. The deflection direction can be reversed by flipping the RS spin (i.e., propagation direction) of the incident plane wave (Supplementary Note V). We conduct microwave experiments to verify the theory by using the experiment setup and metasurface in Fig. 3e (Methods). The measured far-field intensity scattered by the metasurface is shown as the red line in Fig. 3f, which agrees well with the simulation result denoted by the black line. Only the backward scattering lobe exhibits a deflection angle, confirming the validity of the theory.
@@
The synergy of RS and PB phases can enable high-dimensional wavefront manipulation. To demonstrate this, we consider the higher-order RS metasurface in Fig. 4a under the incidence of a plane wave with circularly polarized electric and magnetic fields. Compared to the metasurface in Fig. 3a, the meta-atoms here are rotated around the local $z$ axis, in addition to the rotation around the local $x$ or $y$ axis (Methods). The $z$-axis rotation induces electric and magnetic polarization evolutions, giving rise to the PB phase $\Phi_{\mathrm{pb}}$. Consequently, the metasurface generates scattering fields carrying the total geometric phase $\Phi_{\text{tot }}=\Phi_{\mathrm{pb}}+\Phi_{\mathrm{rs}}$. The incident wave field is $\bm{\Psi}_{\text{in}}=[\left(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\text{in}}\hat{\mathbf{e}}_{y}\right)-\mathrm{i}(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\text{in}}\hat{\mathbf{h}}_{x})]Ee^{-\mathrm{i}k_{0}z-\mathrm{i}\omega t}$, which carries the spins $\sigma_{\text{in }}=-1$ and $\kappa_{\text{in }}=-1$. Its electric, magnetic, and RS polarization states are represented by the south poles on the E-sphere, H-sphere, and RS-sphere, respectively, as shown in Fig. 4b. The incident wave excites the 4D dipole in the meta-atoms: $\mathbf{D}=\mathbf{p}+\mathrm{i}{\mathbf{m}}=[(p_{x}\hat{\mathbf{e}}_{x}+p_{y}\hat{\mathbf{e}}_{y})+{\mathrm{i}}(m_{y}\hat{\mathbf{h}}_{y}+m_{x}\hat{\mathbf{h}}_{x})]e^{-\mathrm{i}\omega t}$. The polarization states of the electric, magnetic, and RS dipoles are represented by $|p\rangle,|m\rangle$, and $\left|D_{\mathrm{rs}}\right\rangle$ on the equators with the azimuthal angles $2\alpha_{\mathrm{p}},2\alpha_{\mathrm{m}}$, and $2\alpha_{\mathrm{rs}}$, respectively. Here, $\alpha_{\mathrm{p}}\left(\alpha_{\mathrm{m}}\right)$ is the orientation angle of the electric (magnetic) dipole in the coordinate frame; $\alpha_{\mathrm{rs}}$ is the orientation angle of the RS dipole in the constitutive frame. The meta-atoms in Fig. 4a are designed to satisfy $\alpha_{\mathrm{p}}=\alpha_{\mathrm{rs}}$. The metasurface generates the scattering field $\bm{\Psi}_{\text{out }}=[(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\text{out}}\hat{\mathbf{e}}_{y})+\mathrm{i}\kappa_{\text{out}}(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\text{out}}\hat{\mathbf{h}}_{x})]Ee^{\mathrm{i}\kappa_{\text{out}}k_{0}z-\mathrm{i}\omega t}$, which has four polarization components $\left|\sigma_{\text{out}}=\pm 1,\kappa_{\text{out}}=\pm 1\right\rangle$, as shown in Fig. 4a. Correspondingly, there are four polarization evolution channels $\left|\sigma_{\mathrm{in}},\kappa_{\mathrm{in}}\right\rangle\rightarrow\left|\sigma_{\text{out}},\kappa_{\text{out}}\right\rangle$. Figure 4b shows the 4D polarization evolution in the channel $|-1,-1\rangle\rightarrow|+1,+1\rangle$ as an example. The resulting geometric phase is:
@@
Figure 4c shows the phases and amplitudes of the four output waves for different orientation angle $\alpha_{\mathrm{rs}}$. The simulated phases (solid lines) are consistent with the total geometric phases (symbols) given by the solid angles on the Poincaré hypersphere. The simulated amplitudes of different outputs are approximately equal and nearly independent of $\alpha_{\mathrm{rs}}$.
@@
Figure 4d shows the simulated electric fields of the four output waves. The wavefronts are consistent with the deflection directions (yellow arrows) predicted based on the geometric phase gradient. Specifically, the output $|+1,+1\rangle$ exhibits the largest deflection angle; $|+1,-1\rangle$ and $|-1,+1\rangle$ exhibit the same deflection angle; $|-1,-1\rangle$ undergoes no deflection. The number of distinct output wavefronts can be increased by flipping the electric spin and RS spin of the incident wave (Supplementary Notes VI and VII). Figure 4e shows the fabricated metasurface sample comprising two supercells. The scattering far-field intensities under the incidence of the plane waves $\left|\sigma_{\mathrm{in}}=\pm 1,\kappa_{\mathrm{in}}=-1\right\rangle$ are shown in Fig. 4f. We notice that a total of eight output intensity lobes (four lobes in each case) emerge in six distinct directions, with consistency between the experimental (red lines) and simulation (black lines) results. Unlike the case in Fig. 4c, the intensities of different outputs are not equal due to the finite size of the metasurface.

### v2

@@ Introduction
The RS phase can emerge in light transmission and reflection at an interface due to 4D polarization evolution, which is determined by the eigen polarization states of the two media forming the interface. Remarkably, the RS phase reveals the hidden geometric nature of the phase shifts governed by Fresnel equations, a cornerstone of classical optics. As shown in Fig. 2a, we consider that a plane wave propagates along +z+z direction and normally impinges on the surface of an isotropic medium with permittivity $\varepsilon$ and permeability $\mu$, giving rise to RS polarization evolution. The incident electric (magnetic) field is linearly polarized in $x$ ($y$) direction. We denote the RS polarization states of the incident, reflected, and transmitted waves as $\left|F_{1}^{\text{in}}\right\rangle,\left|F_{1}^{\text{ref}}\right\rangle$, and $\left|F_{1}^{\text{tra}}\right\rangle$, respectively. The transmission involves the polarization evolution $\left|F_{1}^{\text{in}}\right\rangle\rightarrow\left|F_{1}^{\text{tra}}\right\rangle$, while the reflection involves the polarization evolution $\left|F_{1}^{\text{in}}\right\rangle\rightarrow\left|F_{1}^{\text{tra}}\right\rangle\rightarrow\left|F_{1}^{\text{ref}}\right\rangle$. These RS polarization evolutions induce the RS phases $\Phi_{\mathrm{rs}}^{\text{tra}}$ and $\Phi_{\mathrm{rs}}^{\mathrm{ref}}$, which manifest in the transmitted and reflected waves, respectively. As an example, we assume the medium is a lossless metal. In this case, $\left|F_{1}^{\mathrm{in}}\right\rangle=\frac{1}{\sqrt{2}}(\hat{\mathbf{e}}_{x}+\mathrm{i}\hat{\mathbf{h}}_{y}),\left|F_{1}^{\mathrm{ref}}\right\rangle=\frac{1}{\sqrt{2}}(\hat{\mathbf{e}}_{x}-\mathrm{i}\hat{\mathbf{h}}_{y})$, and $\left|F_{1}^{\text{tra}}\right\rangle=\frac{\sqrt{\left|{\mu}\right|}}{\sqrt{\left|{\mu}\right|+\left|{\varepsilon}\right|}}(\hat{\mathbf{e}}_{x}+\frac{\mathrm{i\sqrt{\varepsilon}}}{\sqrt{\mu}}\hat{\mathbf{h}}_{y})$. Notably, $\left|F_{1}^{\text{tra}}\right\rangle$ is linearly polarized in the constitutive frame with orientation angle $\alpha_{\mathrm{rs}}=\text{arctan}(\frac{\mathrm{i\sqrt{\varepsilon}}}{\sqrt{\mu}})$, which locates on the equator of the RS-sphere with azimuthal angle $\phi_{\mathrm{rs}}=2\alpha_{\mathrm{rs}}$. Thus, changing the material properties $(\varepsilon,\mu)$ leads to rotation of local constitutive frame and thus the variation of $\phi_{\mathrm{rs}}$. Figure 2b shows the polarization evolution paths on the RS-sphere for the transmission and reflection. We calculate the RS phases $\Phi_{\mathrm{rs}}^{\text{tra}}$ and $\Phi_{\mathrm{rs}}^{\text{ref}}$ by evaluating the solid angles $\Omega$ enclosed by the paths. The results are shown as the circles in Fig. 2c for different orientation angle $\alpha_{\mathrm{rs}}$, which agree with the transmission and reflection phases (solid lines) predicted by the Fresnel equations [RN1]. This relationship holds for normal incidence at general interfaces between isotropic media (Supplementary Note IV).
@@
Figure 3: RS geometric phase at a metasurface. (a) Schematic of the RS metasurface under the normally incident plane wave with linearly polarized electric field. The inset shows the structure of the meta-atom. (b) RS polarization evolution on the RS-sphere induced by the metasurface. (c) Theoretical RS geometric phase (symbols), simulated phase (solid lines), and simulated amplitude for the forward and backward scattering fields as a function of the orientation angle $\alpha_{\mathrm{rs}}$ of the RS dipole. (d) Simulated electric field scattered by the RS metasurface. (e) Experiment setup and the fabricated metasurface prototype. (f) Simulated and experimentally measured far-field intensity pattern. The shaded area marks the measurement blind zone [-15 deg, 15 deg] due to the source antenna obstruction.
Both the PB and RS phases can emerge at an interface involving anisotropic media. We consider the interface between air and a non-magnetic medium that has anisotropic in-plane permittivity $\overleftrightarrow{\varepsilon}$ with elements $\varepsilon_{xx}=\varepsilon+\delta\cos\phi,\varepsilon_{yy}=\varepsilon-\delta\cos\phi,\varepsilon_{xy}=\delta\sin\phi e^{-\mathrm{i}\theta}$, and $\varepsilon_{yx}=\delta\sin\phi e^{\mathrm{i}\theta}$. The medium supports two orthogonal 4D eigen polarization states. As shown in Fig. 2d, under the normal incidence of a plane wave with polarization $\left|\Psi_{\text{in}}\right\rangle=\frac{1}{2}[(\hat{\mathbf{e}}_{x}+\mathrm{i}\hat{\mathbf{e}}_{y})+\mathrm{i}(\hat{\mathbf{h}}_{y}-\mathrm{i}\hat{\mathbf{h}}_{x})]$, two transmitted waves $\left|\Psi_{\text{tra}}^{\pm}\right\rangle=\frac{1}{\sqrt{1+|\varepsilon\pm\delta|}}(\mathbf{E}_{\pm}+\mathrm{i}\sqrt{\varepsilon\pm\delta}\mathbf{H}_{\pm})$ and two reflected waves $\left|\Psi_{\text{ref}}^{\pm}\right\rangle=\frac{1}{\sqrt{2}}\left(\mathbf{E}_{\pm}-\mathrm{i}\mathbf{H}_{\pm}\right)$ emerge, where $\mathbf{E}_{+}=\cos\frac{\phi}{2}\hat{\mathbf{e}}_{x}+\sin\frac{\phi}{2}e^{\mathrm{i}\theta}\hat{\mathbf{e}}_{y},\mathbf{E}_{-}=\sin\frac{\phi}{2}e^{-\mathrm{i}\theta}\hat{\mathbf{e}}_{x}-\cos\frac{\phi}{2}\hat{\mathbf{e}}_{y},\mathbf{H}_{+}=\cos\frac{\phi}{2}\hat{\mathbf{h}}_{y}-\sin\frac{\phi}{2}e^{\mathrm{i}\theta}\hat{\mathbf{h}}_{x}$, and $\mathbf{H}_{-}=\sin\frac{\phi}{2}e^{-\mathrm{i}\theta}\hat{\mathbf{h}}_{y}+\cos\frac{\phi}{2}\hat{\mathbf{h}}_{x}$. Correspondingly, four polarization evolutions occur at the interface: $\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{+}\right\rangle,\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{-}\right\rangle,\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{+}\right\rangle\rightarrow\left|\Psi_{\text{ref}}^{+}\right\rangle$, and $\left|\Psi_{\text{in}}\right\rangle\rightarrow\left|\Psi_{\text{tra}}^{-}\right\rangle\rightarrow\left|\Psi_{\text{ref}}^{-}\right\rangle$. As an example, Fig. 2e depicts the polarization evolutions associated with $\left|\Psi_{\text{tra}}^{+}\right\rangle$ and $\left|\Psi_{\text{ref}}^{+}\right\rangle$ for $\varepsilon=1+\mathrm{i},\delta=2,\phi=30^{\circ}$, and $\theta=0^{\circ}$. Notably, the electric and magnetic polarizations evolve along the same pathway on the E-sphere and H-sphere, which are independent of the RS polarization evolution on the RS-sphere. Figure 2f shows the total geometric phases (i.e., sum of PB and RS phases) induced by the four polarization evolutions for different $\alpha_{\mathrm{e}}=\frac{\phi}{2}$. The geometric phases (circles) agree with the transmission and reflection phases (solid lines) predicted by the Fresnel equations. Thus, the well-known reflection and transmission phases at interfaces can be interpretated as the geometric phases induced by 4D polarization evolution, uncovering the geometric origin of this fundamental optical phenomena.
RS phase at metasurfaces
@@
Figure 4: Complementary RS and PB phases at a metasurface. (a) Schematic of the quadruplex RS meta-deflector under the normal incidence of a plane wave with LCP electric field. (b) 4D polarization evolution on the Poincaré hypersphere induced by the higher-order RS metasurface. (c) Theoretical geometric phase (symbols), simulated phase (solid lines), and simulated amplitude in four polarization evolution channels as a function of the orientation angle of RS dipole in the meta-atom. (d) Simulated electric field distribution of four output waves. (e) Fabricated metasurface prototype. (f) Simulated and experimentally measured far-field intensity pattern under orthogonal incidence. The shaded area marks the measurement blind zone [-15 deg, 15 deg] due to the source antenna obstruction.
Figure 3b shows the RS polarization evolution in the channel $|-1\rangle\rightarrow|+1\rangle$. The incident state $\left|F_{1}^{\text{in}}\right\rangle$ and backward scattering state $\left|F_{1}^{\text{out}}\right\rangle$ are represented by the south and north poles, respectively. The polarization state of the RS dipole $\mathbf{D}_{\mathrm{rs}}$, denoted as $\left|D_{\mathrm{rs}}\right\rangle$, is located on the equator. The orientation angle of $\mathbf{D}_{\mathrm{rs}}$ in the constitutive frame is $\alpha_{\mathrm{rs}}=\text{arctan}(\frac{\mathrm{i}m_{y}}{p_{x}})$. Rotating the meta-atoms around $x$ or $y$ direction will alter $\alpha_{\mathrm{rs}}$ (Supplementary Note V) and change the azimuthal angle on the RS-sphere. The resulting RS phase is $\Phi_{\mathrm{rs}}=-2\alpha_{\mathrm{rs}}$. Notably, the PB phase vanishes in this case because the meta-atom rotation around $x$ or $y$ direction does not change the electric or magnetic polarization. Figure 3c shows the phases and amplitudes of the scattering fields as a function of the orientation angle $\alpha_{\mathrm{rs}}$ of the RS dipole $\mathbf{D}_{\mathrm{rs}}$. The simulated phases (solid lines) agree with the RS phase (circles) obtained by evaluating the solid angle in the RS-sphere. Notably, the phase of the backward scattering field $|+1\rangle$ exhibits a linear relationship with $\alpha_{\mathrm{rs}}$, while the phase of the forward scattering field $|-1\rangle$ is independent of $\alpha_{\mathrm{rs}}$. In addition, the scattering amplitudes remain approximately constant for different $\alpha_{\mathrm{rs}}$.
@@
The RS metasurface in Fig. 3a can generate RS phase gradient along $y$ direction, which can deflect the incident wave. Figure 3d shows the simulated electric field $E_{x}$ scattered by the RS metasurface. We note that only the backward scattering field is deflected into oblique direction, and the wavefront is consistent with the deflection angle $\beta=\arcsin(\frac{1}{k_{0}}\frac{\Delta\Phi_{\mathrm{rs}}}{\Delta y})$ predicted by the generalized Snell’s law [RN59]. The deflection direction can be reversed by flipping the RS spin (i.e., propagation direction) of the incident plane wave (Supplementary Note VI). We conduct microwave experiments to verify the theory by using the experiment setup and metasurface in Fig. 3e (Methods). The measured far-field intensity scattered by the metasurface is shown as the red line in Fig. 3f, which agrees well with the simulation result denoted by the black line. Only the backward scattering lobe exhibits a deflection angle, confirming the validity of the theory.
@@
To verify the high-dimensional geometric framework based on the RS and PB phases, we consider the higher-order RS metasurface in Fig. 4a under the incidence of a plane wave with circularly polarized electric and magnetic fields. Compared to the metasurface in Fig. 3a, the meta-atoms here are further rotated around the local $z$ axis (Methods), which induces electric and magnetic polarization evolutions and gives rise to the additional PB phase $\Phi_{\mathrm{pb}}$. Consequently, the metasurface generates scattering fields carrying the total geometric phase $\Phi_{\text{tot }}=\Phi_{\mathrm{pb}}+\Phi_{\mathrm{rs}}$. The incident wave field is $\bm{\Psi}_{\text{in}}=[\left(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\text{in}}\hat{\mathbf{e}}_{y}\right)-\mathrm{i}(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\text{in}}\hat{\mathbf{h}}_{x})]Ee^{-\mathrm{i}k_{0}z-\mathrm{i}\omega t}$, which carries the spins $\sigma_{\text{in }}=-1$ and $\kappa_{\text{in }}=-1$. Its electric, magnetic, and RS polarization states are represented by the south poles on the E-sphere, H-sphere, and RS-sphere, respectively, as shown in Fig. 4b. The incident wave excites the 4D dipole in the meta-atoms: $\mathbf{D}=\mathbf{p}+\mathrm{i}{\mathbf{m}}=[(p_{x}\hat{\mathbf{e}}_{x}+p_{y}\hat{\mathbf{e}}_{y})+{\mathrm{i}}(m_{y}\hat{\mathbf{h}}_{y}+m_{x}\hat{\mathbf{h}}_{x})]e^{-\mathrm{i}\omega t}$. The polarization states of the electric, magnetic, and RS dipoles are represented by $|p\rangle,|m\rangle$, and $\left|D_{\mathrm{rs}}\right\rangle$ on the equators with the azimuthal angles $2\alpha_{\mathrm{p}},2\alpha_{\mathrm{m}}$, and $2\alpha_{\mathrm{rs}}$, respectively. Here, $\alpha_{\mathrm{p}}\left(\alpha_{\mathrm{m}}\right)$ is the orientation angle of the electric (magnetic) dipole in the coordinate frame; $\alpha_{\mathrm{rs}}$ is the orientation angle of the RS dipole in the constitutive frame. The meta-atoms in Fig. 4a are designed to satisfy $\alpha_{\mathrm{p}}=\alpha_{\mathrm{rs}}$. The metasurface generates the scattering field $\bm{\Psi}_{\text{out }}=[(\hat{\mathbf{e}}_{x}+\mathrm{i}\sigma_{\text{out}}\hat{\mathbf{e}}_{y})+\mathrm{i}\kappa_{\text{out}}(\hat{\mathbf{h}}_{y}-\mathrm{i}\sigma_{\text{out}}\hat{\mathbf{h}}_{x})]Ee^{\mathrm{i}\kappa_{\text{out}}k_{0}z-\mathrm{i}\omega t}$, which has four polarization components $\left|\sigma_{\text{out}}=\pm 1,\kappa_{\text{out}}=\pm 1\right\rangle$, as shown in Fig. 4a. Correspondingly, there are four polarization evolution channels $\left|\sigma_{\mathrm{in}},\kappa_{\mathrm{in}}\right\rangle\rightarrow\left|\sigma_{\text{out}},\kappa_{\text{out}}\right\rangle$. Figure 4b shows the 4D polarization evolution in the channel $|-1,-1\rangle\rightarrow|+1,+1\rangle$ as an example. The resulting geometric phase is:
@@
Figure 4c shows the phases and amplitudes of the four output waves for different orientation angle $\alpha_{\mathrm{rs}}$. The simulated phases (solid lines) are consistent with the total geometric phases (symbols) given by the solid angles on the Poincaré hypersphere, confirming the emergence of both the PB and RS phases. The simulated amplitudes of different outputs are approximately equal and nearly independent of $\alpha_{\mathrm{rs}}$.
@@
Figure 4d shows the simulated electric fields of the four output waves. The wavefronts are consistent with the deflection directions (yellow arrows) predicted based on the geometric phase gradient. Specifically, the output $|+1,+1\rangle$ exhibits the largest deflection angle; $|+1,-1\rangle$ and $|-1,+1\rangle$ exhibit the same deflection angle; $|-1,-1\rangle$ undergoes no deflection. The number of distinct output wavefronts can be increased by flipping the electric spin and RS spin of the incident wave (Supplementary Notes VII and VIII). Figure 4e shows the fabricated metasurface sample comprising two supercells. The scattering far-field intensities under the incidence of the plane waves $\left|\sigma_{\mathrm{in}}=\pm 1,\kappa_{\mathrm{in}}=-1\right\rangle$ are shown in Fig. 4f. We notice that a total of eight output intensity lobes (four lobes in each case) emerge in six distinct directions, with consistency between the experimental (red lines) and simulation (black lines) results. Unlike the case in Fig. 4c, the intensities of different outputs are not equal due to the finite size of the metasurface.